OpenStudy (anonymous):

Some antibiotics can harm the ribosomes in animal cells. Which of the following would you predict to be the most likely long-term effect of damage to ribosomes? A. Substances would not be able to enter or leave the cell, and affected cells would eventually burst. B. Proteins would not be modified after being synthesized, and would be unable to function properly. C. Cells would die because they would not be able to create the proteins needed to carry out life functions. D. Errors would be made during DNA replication, and cells would then be unable to grow and divide.

OpenStudy (anonymous):

@britneyspears

OpenStudy (anonymous):

b. Proteins would not be modified after being synthesized, and would be unable to function properly.

OpenStudy (anonymous):

YOUR THE BEST

OpenStudy (anonymous):

Thank you!! Xo!

OpenStudy (anonymous):

Can you help with another?

OpenStudy (anonymous):

@britneyspears

OpenStudy (anonymous):

Sure!! :)

OpenStudy (anonymous):

What characteristic do the nucleus, the mitochondria, and the vacuoles of animal cells have in common? A. They all create energy while doing their tasks. B.They all contain some of the cell's DNA. C. They are all enclosed by membranes made of phospholipids. D. They are all involved in producing food for the cell.

OpenStudy (anonymous):

c. They are all enclosed by membranes made of phospholipids.

OpenStudy (anonymous):

Thank you..Hb another?

OpenStudy (anonymous):

Sure!! Xo! :)

OpenStudy (anonymous):

Cell membranes are structured so that molecules can pass in and out of the cell across them. While both plant and animal cells have membranes, plant cells also have cell walls surrounding their cells. Which of the following observations BEST supports the conclusion that molecules can cross cell walls as well? Every cell has its own nucleus. The chloroplasts need light in order to work. The cell wall is directly outside the cell membrane. Water from the soil is used in photosynthesis. @britneyspears

OpenStudy (anonymous):

Water from the soil is used in photosynthesis.

OpenStudy (anonymous):

@britneyspears you're helping me out so much see im in 9th grade and I have a daughter already and it's so hard to do school when she is crying and ect being a 1 year old lol I had her when I was 15 im 16 now so u are a great help

OpenStudy (anonymous):

Aww, thank you!! Xo!

OpenStudy (anonymous):

Can you help me with a few more?

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

If an animal cell has excess energy, what is most likely to happen to this energy? It will be considered waste. It will be stored in high energy phosphate bonds. It will be converted into starch. It will be stored as fat molecules.

OpenStudy (anonymous):

It will be stored as fat molecules.

OpenStudy (anonymous):

If energy is needed to remove a phosphate group from a chain in ATP, you can conclude that the energy needed for production must be Found outside the cell to maintain balance within Proportional to the amount of food taken in by the cell Equal to the amount of energy produced Less than the amount of energy produced

OpenStudy (anonymous):

Less than the amount of energy produced.

OpenStudy (anonymous):

How might an increase in the rate of photosynthesis affect the conditions necessary for cellular respiration? Increased oxygen and glucose from photosynthesis would increase cellular respiration. Increased water and carbon dioxide from photosynthesis would increase cellular respiration. Increased water and carbon dioxide from photosynthesis would increase the amount of convertible ADP. Increased carbon dioxide and glucose from photosynthesis would increase cellular metabolism

OpenStudy (anonymous):

Increased oxygen and glucose from photosynthesis would increase cellular respiration.

OpenStudy (anonymous):

What are the reactants for cellular respiration? H2O and O2 Glucose and CO2 H2O and CO2 Glucose and O2

OpenStudy (anonymous):

H2O and CO2.

OpenStudy (anonymous):

When the carbon atoms of the glucose molecule are broken apart in glycolysis and the Krebs cycle, what is the result? CO2 and ATP C6H12O6 and H2O FADH2 and ATP CO2 and H2O

OpenStudy (anonymous):

CO2 and H2O.

OpenStudy (anonymous):

A plant has to produce carbohydrates because carbohydrates are needed to move ATP from cell to cell. carbohydrates are needed for building material. carbohydrates store the genetic material within the cells. carbohydrates store energy for a long time

OpenStudy (anonymous):

Carbohydrates store energy for a long time.

OpenStudy (anonymous):

Plants growing indoors tend to bend their stems so that their leaves turn toward a window. What does this suggest about the levels of reactants for photosynthesis? The levels of carbon dioxide are too low. The levels of carbohydrates are too low. The levels of oxygen are too low. The levels of light are too low.

OpenStudy (anonymous):

The levels of light are too low.

OpenStudy (anonymous):

You have been a life saver today thank you for everything

OpenStudy (anonymous):

You're welcome!! Xo! Is this all for today?

OpenStudy (anonymous):

All for biology you good at English?

OpenStudy (anonymous):

Maybe, but I can try.

OpenStudy (anonymous):

ok, Why do some readers choose to join reader groups? Groups offer opportunities to discuss interesting topics in literature. Groups allow readers to find out the ends of stories without reading. Groups usually give away free books to their regular attendees. Groups can prove to the community that you are a true “reader.”

OpenStudy (anonymous):

Groups offer opportunities to discuss interesting topics in literature.

OpenStudy (anonymous):

What can diaries and journals do for readers? Diaries are not helpful, but journals can improve comprehension. Journals are not helpful, but diaries can improve comprehension. Neither diaries nor journals are recommended for readers to use. Both diaries and journals help readers keep track of information.

OpenStudy (anonymous):

Both diaries and journals help readers keep track of information.

OpenStudy (anonymous):

That's what I thought. Which of the following is a strong post-reading strategy? Paraphrase each line of the entire story by hand. Write an essay without looking at your notes. Discuss difficult concepts with other people. Loan the book to a friend to read after you.

OpenStudy (anonymous):

Discuss difficult concepts with other people.

OpenStudy (anonymous):

ok, any good at math?

OpenStudy (anonymous):

Truly yeah!! :)

OpenStudy (anonymous):

Yayy. ok Name the algebraic property demonstrated in the example below: (2 + x) + y = 2 + (x + y) Commutative Property of Addition Associative Property of Addition Distributive Property Transitive Property

OpenStudy (anonymous):

Associative Property of Addition. When three or more numbers are added, the sum is the same regardless of the grouping of the addends. For example, (2 + 3) + 4 = 2 + (3 + 4).

OpenStudy (anonymous):

Ok, thank you and next is... Solve 3c + 4 = -20. -5.3 6 8 -8

OpenStudy (anonymous):

Simplifying 3c + 4 = -20 Reorder the terms: 4 + 3c = -20 Solving 4 + 3c = -20 Solving for variable 'c'. Move all terms containing c to the left, all other terms to the right. Add '-4' to each side of the equation. 4 + -4 + 3c = -20 + -4 Combine like terms: 4 + -4 = 0 0 + 3c = -20 + -4 3c = -20 + -4 Combine like terms: -20 + -4 = -24 3c = -24 Divide each side by '3'. c = -8 Simplifying c = -8

OpenStudy (anonymous):

Thank you as always... A water park charges $5 for entry into the park and an additional $2 for each of the big waterslides. Steven spent $17 on his visit to the water park. Which equation can be used to find the number of big waterslides, x, that Steven went on? 5 + 2x = 17 5x + 2 = 17 5x + 2x = 17 5 + 2 = 17x

OpenStudy (anonymous):

a. 5 + 2x = 17 Because, 17 - 5 = 12. 12 divided by 2 = 6. He went on 6 big waterslides.

OpenStudy (anonymous):

Solve for x: (x+4) / 2 = 7

OpenStudy (anonymous):

(x + 4) / 2 = 7 Multiply entire equation by 2 to get rid of fraction. (x + 4) = 14 Remove parentheses. x + 4 = 14 Subtract 4 from both sides. x = 10

OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below: 4(x + y) = 4x + 4y Distributive Property Transitive Property Associative Property of Multiplication Commutative Property of Multiplication

OpenStudy (anonymous):

Distributive Property. The Distributive Property states that the product of a number and a sum is equal to the sum of the individual products of the addends and the number. That is, a(b + c) = ab + ac.

OpenStudy (anonymous):

Solve for x: 1 over 4 (2x − 14) = 4 15 15 over 2 9 11

OpenStudy (anonymous):

1/4 (2x - 14) = 4 4 * (1/4) (2x - 14) = 4 * 4 2x - 14 = 16 2x -14 + 14 = 16 + 14 2x = 30 x = 15

OpenStudy (anonymous):

Solve y over negative 6 + 5 = 9. 24 -24 78 -78

OpenStudy (anonymous):

The solution is: y = -24.

OpenStudy (anonymous):

Solve 4(x + 1) + 8 = 24. 3.75 -2 5 3

OpenStudy (anonymous):

Distribute. 4x + 4 + 8 = 24. Add like terms. 4x + 12 = 24. Subtract 12 both sides. 4x = 12. Divide. x = 3.

OpenStudy (anonymous):

Solve 4y - y + 11 = 38. 9 -9 3.8 8.6

OpenStudy (anonymous):

4y - y + 11 = 38 3y = 38 - 11 3y = 27 y = 27 / 3 y = 9

OpenStudy (anonymous):

Which of the following describes the correct process for solving the equation 2x - 6 = 22 and arrives at the correct solution? Divide both sides by -6 and then add 2. The solution is x = 8. Subtract 6 from both sides of the equation and then divide by 2. The solution is x = 7. Add 6 to both sides of the equation and then divide by 2. The solution is x = 14. Add 6 to both sides of the equation and then divide by 22. The solution is x = -14.

OpenStudy (anonymous):

Add 6 to both sides of the equation and then divide by 2. The solution is x = 14.

OpenStudy (anonymous):

Can you help me with two more English questions?

OpenStudy (anonymous):

Yeah :)

OpenStudy (anonymous):

Great :) Create a paraphrase of the following sentence from “The Cold Equations”: There could be no alternative—but it required a few moments of conditioning for even an EDS pilot to prepare himself to walk across the room and coldly, deliberately, take the life of a man he had yet to meet.

OpenStudy (anonymous):

Honestly, I cannot do this one, sorry.

OpenStudy (anonymous):

Its ok thank thoo the other one is about the same thing so ill try to figure it out thanks for everything today?

OpenStudy (anonymous):

You're welcome!! And If you need more help, then come by me anytime!! Xo!

OpenStudy (anonymous):

Will do thanks :)